A complete demand function for Titan Watch is given by the following equation: - 𝑄𝑑 = βˆ’30𝑃 + 0.05π‘Œ + 2 Pr + 4𝑇 Where, P=Price of Titan Watch, Y=Income, Pr= Price of related goods, T= Taste and 𝑄𝑑= Quantity demand for watch. (i) Draw demand curve assuming Y= 5000, Pr = 25 and T=30. (ii) What will happen if T increases to 40? Show it graphically


1
Expert's answer
2021-07-07T08:54:49-0400

(i) If Y = 5000, Pr = 25 and T=30, then the demand function is:

"Qd = \u221230\ud835\udc43 + 0.05\u00d75,000 + 2\u00d725 + 4\u00d730 = 420 - 30P," so the demand curve is a downward-sloping line.

(ii) If T increases to 40, then:

Qd = 460 - 30P, so the demand will increase and the demand curve will shift upward.